# Tropical Gradient Descent

### v2

@@ Abstract
We propose a gradient descent method for solving optimization problems arising in settings of tropical geometry—a variant of algebraic geometry that has become increasingly studied in applications such as computational biology, economics, and computer science. Our approach takes advantage of the polyhedral and combinatorial structures arising in tropical geometry to propose a versatile approach for approximating local minima in tropical statistical optimization problems—a rapidly growing body of work in recent years. Theoretical results establish global solvability for 1-sample problems and a convergence rate matching classical gradient descent. Numerical experiments demonstrate the method’s superior performance over classical descent for tropical optimization problems which exhibit tropical convexity but not classical convexity. We also demonstrate the seamless integration of tropical descent into advanced optimization methods, such as Adam, offering improved overall performance.
@@ 1 Introduction
Where algebraic geometry studies geometric properties of solution sets of systems of multivariate polynomials, tropical geometry restricts to the case of polynomials defined by linearizing operations, where the “sum” of two elements is their maximum and the “product” of two elements is their sum. Evaluating polynomials with these operations results in piecewise linear functions. Tropical geometry is a relatively young field of pure mathematics established in the 1990s (with roots dating back further) and has recently become an area of active interest in computational and applied mathematics for its relevance to the statistical analyses of phylogenetic trees (Monod et al.,, 2018), the problem of dynamic programming in computer science (Maclagan & Sturmfels,, 2015), and in mechanism design for two-player games in game theory (Lin & Tran,, 2019). Following the identification of the tropical Grassmannian and the space of phylogenetic trees (Speyer & Sturmfels,, 2004) in particular, there has been surge of research (Barnhill et al.,, 2024b; Lin & Yoshida,, 2018; Comăneci & Joswig,, 2024; Monod et al.,, 2018) which necessitates the development of new computational techniques for optimisation in tropical geometric settings.
@@
The training of machine learning models and the vast majority of computational tasks in statistics entail solving optimization problems; this paper is motivated by statistical optimization problems in tropical geometric settings. Initial research into tropical statistics considered optimization problems such as the identification of Fermat–Weber points which, due to the piecewise linear nature of tropical geometry, can be re-framed as a linear program (Comăneci & Joswig,, 2024; Lin & Yoshida,, 2018). Unfortunately, the complexity of these linear programs often scales with the size of our dataset, so gradient descent is more commonly implemented in practice (Barnhill et al.,, 2024a, b). More recently—following the identification of ReLU neural networks and tropical rational functions (Zhang et al.,, 2018)—more sophisticated neural network optimization problems are being studied in the tropical setting (Pasque et al.,, 2024; Yoshida et al.,, 2024), which necessitates the use of gradient methods for optimization in the tropical setting. In this paper, we address the natural question—given the need for gradient methods in tropical data science, how can we tailor classical gradient methods to the tropical setting? We provide a foundational tropical gradient descent framework to solve the optimization problems which arise in the computation of a wide range of tropical statistics, an area of active interest in applied tropical geometry.
@@
The remainder of this paper is organized as follows. The following section contains the necessary foundations of tropical geometry, an introduction to tropical location problems (Comăneci,, 2024), and the array of statistical optimization tasks studied in this paper. Section 3 contains the theoretical contributions of this paper; we formulate steepest descent with respect to the tropical norm and present the theoretical guarantees for tropical descent, most notably its convergence for a wide class of tropically quasi-convex functions and a convergence rate of $O(1/\sqrt{k})$. In Section 4, we perform a comprehensive numerical study of our proposed tropical descent method and its performance in the statistical optimization problems of interest. We conclude with a discussion of possible future work on tropical gradient methods in Section 5.
@@ 2 Background and Preliminaries
In this section, we overview the background on tropical geometry and present the statistical optimization problems which we study in the tropical setting.
@@ 2.1 Tropical Geometry
Here, we review the algebraic and geometric structure of the tropical projective torus, the state space for phylogenetic data and the domain of our tropical statistical optimization problems.
@@
The additive identity is $\infty$- ∞ and the multiplicative identity is 00. Tropical subtraction is not defined; tropical division is given by classical subtraction.
The tropical algebra defined above is sometimes referred to as the max-plus algebra; we can define the min-plus algebra analogously, taking $\min\{a,b\}=:a\oplus b$ as the additive operation on ${}\mathbb{R}\cup\{\infty\}$. These are algebraically equivalent through negation. Unless specified, we use the max-plus convention as this is better suited for phylogenetic statistics on the tropical projective torus (Speyer & Sturmfels,, 2004).
@@
The $n$-dimensional tropical projective torus is a quotient space constructed by endowing $\mathbb{R}^{N}$ with the equivalence relation
$\mathbf{x}\sim\mathbf{y}\Leftrightarrow\exists\,a:\>\mathbf{x}=a\odot\mathbf{y};$ (1)
@@
While a single point in the tropical projective torus is given by some ray $\mathbf{x}+\mathbb{R}\mathbf{1}$, we note that there is a unique representative of this equivalence class whose coordinates sum to zero. In taking such representatives, we can identify the tropical projective torus $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ and the hyperplane =$\mathcal{H}=\{\mathbf{x}\in\mathbb{R}^{N}:\sum x_{i}=0\}$. Throughout this paper, we use this homeomorphism to visualize the tropical projective torus (e.g., Figures 1 and 2).
The tropical projective torus is the ambient space containing the tropical Grassmannian, which is equivalent to the space of phylogenetic trees (Speyer & Sturmfels,, 2004). This has sparked the study of tropical data science, formalizing statistical techniques which respect the tropical geometry of the state space (Monod et al.,, 2018; Yoshida,, 2022).
The generalized Hilbert projective metric defined above is widely accepted and preferred for the development of geometric statistical tools on the tropical projective torus (Yoshida,, 2022). However, recent work has shown theoretical advantages to the use of an asymmetric metric (Comăneci & Joswig,, 2024).
@@
Figure 1: The $d_{\mathrm{tr}}$, $d_{\triangle_{\min}}$ and $d_{\triangle_{\max}}$ tropical unit balls in $\mathbb{R}^{3}/\mathbb{R}\mathbf{1}\cong\mathcal{H}=\{\sum x_{i}=0\}$.
@@
$\mathcal{H}_{\mathbf{a}}=\{\mathbf{x}:\exists\,i\neq j\text{ s.t. }a_{i}+x_{i}=a_{j}+x_{j}=\max\{a_{k}+x_{k}\}\}.$
@@
Definition 8 (Tropical Convex Hull (Develin & Sturmfels,, 2004)).
A set $S\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ is $\max$-tropically convex if it contains $\alpha\odot\mathbf{x}\boxplus\beta\odot\mathbf{y}$ for all $\mathbf{x},\mathbf{y}\in S$ and $\alpha,\beta\in\mathbb{R}$. For a finite subset $X=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\}\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, the max-plus tropical convex hull of $X$ is the set of all max-tropical linear combinations of points in $X$,
@@ 2.2 Tropical Location Problems
Tropical location problems are a new avenue of research, motivated by the optimization of statistical loss functions; they involve the minimization over some dataset of some loss function which heuristically increases with distance from some kernel. Here we outline the relevant definitions of tropical location problems and their theoretical behavior as presented by Comăneci, (2024). We introduce the problem of tropical linear regression (Akian et al.,, 2023) as motivation.
@@ 2.2.1 Motivating Example
The problem of tropical linear regression looks to find the best-fit tropical hyperplane which minimizes the maximal distance to a set of data points. While this is an inherently statistical optimization problem, it has recently been proven to be polynomial-time equivalent to solving mean payoff games (Akian et al.,, 2023).
A tropical hyperplane is uniquely defined by a cone point $\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, allowing us to formulate the best-fit tropical hyperplane as an optimization problem over $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$.
Definition 9 (Tropical Linear Regression (Akian et al.,, 2023)).
@@
$\min_{\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}}f(\mathbf{t})=\max_{k\leq K}d_{\mathrm{tr}}(\mathbf{x},\mathcal{H}_{\mathbf{t}})=\max_{k\leq K}(\mathbf{x}_{k}-\mathbf{t})_{i}-(\mathbf{x}_{k}-\mathbf{t})_{j}$
where $i,j$ are such that $\forall\,\ell\neq i,j:(\mathbf{x}_{k}-\mathbf{t})_{i}\geq(\mathbf{x}_{k}-\mathbf{t})_{j}\geq(\mathbf{x}_{k}-\mathbf{t})_{\ell}$.
In Figure 3, we see a heat map of this objective function over the tropical projective plane for a sample of size 3. We see immediately that the objective is not convex. The linear regression objective has sparse gradient almost everywhere, which produces valleys. These properties are considered unfavourable extreme cases for gradient methods, but are particularly common in tropical optimization problems. While the tropical linear regression problem is not convex, it is quasi-convex along tropical line segments; that is, its sub-level sets are tropically convex.
@@ 2.2.2 Quasi-Convexity of Tropical Functions
In formalizing tropical location problems, the work of Comăneci, (2024) first considers functions which increase along tropical geodesics from a kernel $\mathbf{v}$; that is, for any point $\mathbf{t}$ on a geodesic from $\mathbf{v}$ to $\mathbf{w}$, we have that $f(\mathbf{t})\leq f(\mathbf{w})$. This property is equivalent to star-convexity of sublevel sets. To formally state this, we define (oriented) geodesic segments and $\triangle_{\min}$-star convexity of sets.
@@
The (oriented) geodesic segment between $\mathbf{a},\mathbf{b}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, is given by $[\mathbf{a},\mathbf{b}]_{\triangle_{\min}}\coloneqq\{\mathbf{x}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}:d_{\triangle_{\min}}(\mathbf{a},\mathbf{x})+d_{\triangle_{\min}}(\mathbf{x},\mathbf{b})=d_{\triangle_{\min}}(\mathbf{a},\mathbf{b})\}$.
@@
A set $K\subseteq\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ is a $\triangle_{\min}$-star-convex set with kernel $\mathbf{v}$ if, for every $\mathbf{w}\in K$, we have $[\mathbf{v},\mathbf{w}]_{\triangle_{\min}}\subseteq K$.
Using the above definitions, we can characterize a loss function which is increasing along tropical geodesics as a function with $\triangle_{\min}$-star-convex sublevel sets. The following definition gives a closed form expression for such functions, while Theorem 13 proves their equivalence.
Definition 12 ($\triangle_{\min}$-Star-Quasi-Convex Functions (Comăneci,, 2024)).
@@
Theorem 13 (Theorem 17 of (Comăneci,, 2024)).
@@
These $\triangle_{\min}$-star-convex functions act as a loss function with respect to a single data point. For example, the tropical metric $d_{\mathrm{tr}}$ is both $\triangle_{\min}$ and $\triangle_{\max}$-star-quasi-convex. The distance to a hyperplane $d_{\mathrm{tr}}(\mathbf{x},\mathcal{H}_{\mathbf{t}})$ is $\triangle_{\max}$-star-quasi-convex in $\mathbf{x}$, but $\triangle_{\min}$-star-quasi-convex in $\mathbf{t}$.
@@
Proposition 15 (Proposition 8 of (Comăneci,, 2024)).
Any $\triangle_{\min}$-star-convex set is $\min$-tropically convex. Hence, any $\triangle_{\min}$-star-quasi-convex function is tropically quasi-convex in that it has tropically convex sublevel sets.
@@ 2.2.3 Tropical Location Problems
The $\triangle_{\min}$-star-convex functions defined above are generally a measure of closeness to the kernel $\mathbf{v}$, but in minimizing a statistical loss function we look at closeness to a sample of points. A location problem (Laporte et al.,, 2019) measures the closeness to a sample rather than a single kernel.
Definition 16 (Min(/max)-Tropical Location Problem (Comăneci,, 2024)).
Let $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ be our dataset. For $k\leq K$, let $h_{k}$ be a $\triangle_{\min}$-star-quasi-convex function with kernel $\mathbf{x}_{k}$, and let $g:\mathbb{R}^{K}\rightarrow\mathbb{R}$ be increasing in every coordinate. A min-tropical location problem is the minimization of an objective function $f:\mathbb{R}^{N}/\mathbb{R}\mathbf{1}\rightarrow\mathbb{R}$ given by $f=g(h_{1},\dots,h_{K})$. A max-tropical location problem is defined similarly, where the $h_{k}$ are $\triangle_{\max}$-star-quasi-convex.
The main result of (Comăneci,, 2024) is the following: tropical location problems contain a minimum in the tropical convex hull of its dataset $X$.
Theorem 17 (Theorem 20 of (Comăneci,, 2024)).
@@
We can now revisit our motivating example (Definition 9), showing that tropical linear regression is not just a min-tropical location problem, but also demonstrates tropical quasi-convexity in that it has tropically convex sublevel sets.
@@
The tropical linear regression problem is a $\min$-tropical location problem with tropically convex sublevel sets.
@@
$f(\mathbf{t})$ $=\max_{k\leq K}\min_{j}[\max_{i}[x_{ki}-t_{i}]-x_{kj}+t_{j}]$
$=\max_{k\leq K}\min_{j}[t_{j}-x_{kj}-\min_{i}[t_{i}-x_{ki}]]$
@@
$\gamma(\mathbf{t})$ $=\min_{j}t_{i}.$
@@
As the $h_{k}$ are $\triangle_{\min}$-star-convex, their sublevel sets are min-tropically convex by Proposition 15. This is preserved when taking a maximum over samples, so $f$ also has $\min$-tropically convex sublevel sets. ∎

### v3

@@ Abstract
We propose a gradient descent method for solving optimization problems arising in settings of tropical geometry—a variant of algebraic geometry that has attracted growing interest in applications such as computational biology, economics, and computer science. Our approach takes advantage of the polyhedral and combinatorial structures arising in tropical geometry to propose a versatile method for approximating local minima in tropical statistical optimization problems—a rapidly growing body of work in recent years. Theoretical results establish global solvability for 1-sample problems and a convergence rate matching classical gradient descent. Numerical experiments demonstrate the method’s superior performance compared to classical gradient descent for tropical optimization problems which exhibit tropical convexity but not classical convexity. We also demonstrate the seamless integration of tropical descent into advanced optimization methods, such as Adam, offering improved overall accuracy.
@@ 1 Introduction
Where algebraic geometry studies geometric properties of solution sets of systems of multivariate polynomials, tropical geometry restricts to the case of polynomials defined by linearizing operations, where the “sum” of two elements is their maximum and the “product” of two elements is their sum. Evaluating polynomials with these operations results in piecewise linear functions. Tropical geometry is a relatively young field of pure mathematics established in the 1990s (with roots dating back further) and has recently become an area of active interest in computational and applied mathematics for its relevance to the statistical analyses of phylogenetic trees (monod2018StatPerspective), the problem of dynamic programming in computer science (maclagan2015introduction), and in mechanism design for two-player games in game theory (LIN2019133). Following the identification of the tropical Grassmannian and the space of phylogenetic trees (speyer2004grassmannian) in particular, there has been a surge of research (barnhill2023tml; lin2018FermatWeber; comuaneci2023transportation; monod2018StatPerspective) which necessitates the development of new computational techniques for optimization in tropical geometric settings.
@@
Training machine learning models and many computational tasks in statistics entail solving optimization problems; this paper is motivated by statistical optimization problems in tropical geometric settings. Early research in tropical statistics focused on optimization problems such as the identification of Fermat–Weber points which, due to the piecewise linear nature of tropical geometry, can be re-framed as a linear program (comuaneci2023transportation; lin2018FermatWeber). Unfortunately, the complexity of these linear programs often scales poorly with the size of our dataset, so gradient methods are more commonly implemented in practice (barnhill2024polytropes; barnhill2023tml). More recently—following the identification of ReLU neural networks and tropical rational functions (zhang2018tropical)—more sophisticated neural network optimization problems are being studied in the tropical setting (pasque2024tropDBs; yoshida2023tropNNs), which necessitates the use of gradient methods for optimization in the tropical setting. In this paper, we address the natural question—given the need for gradient methods in tropical data science, how can we tailor classical gradient methods to the tropical setting? We provide a foundational tropical gradient descent framework to solve the optimization problems which arise in the computation of a wide range of tropical statistics, an area of active interest in applied tropical geometry.
@@
The remainder of this paper is organized as follows. The following section contains the necessary foundations of tropical geometry, an introduction to tropical location problems (comuaneci2024location), and an overview of the statistical optimization tasks studied in this paper. Section 3 contains the theoretical contributions of this paper; we formulate steepest descent with respect to the tropical norm and present the theoretical guarantees for tropical descent, most notably its convergence for a wide class of tropically quasi-convex functions and a convergence rate of $O(1/\sqrt{m})$. In Section 4, we perform a comprehensive numerical study of our proposed tropical descent method and its performance in the statistical optimization problems of interest. We conclude with a discussion of possible future work on tropical gradient methods in Section 5.
@@ 2 Background and Preliminaries
In this section, we present a summary of tropical geometry and the statistical optimization problems which we study in the tropical setting.
@@ 2.1 Tropical Geometry
We begin by reviewing the algebraic and geometric structure of the tropical projective torus, the state space for phylogenetic data and the domain of our tropical statistical optimization problems.
@@
The additive identity is -$\infty$ and the multiplicative identity is 0. Tropical subtraction is not defined; tropical division is given by classical subtraction.
The tropical algebra defined above is sometimes referred to as the max-plus algebra; we can define the min-plus algebra analogously, taking $\min\{a,b\}=:a\oplus b$ as the additive operation on $\mathbb{R}\cup\{\infty\}$. These are algebraically equivalent under negation. Unless specified, we use the max-plus convention as this is better suited for phylogenetic statistics on the tropical projective torus (speyer2004grassmannian).
@@
The N1N-1-dimensional tropical projective torus is a quotient space constructed by endowing $\mathbb{R}^{N}$ with the equivalence relation
$\mathbf{x}\sim\mathbf{y}\Leftrightarrow\exists\,a{\,\in\mathbb{R}}:\>\mathbf{x}=a\odot\mathbf{y};$ (1)
@@
While a single point in the tropical projective torus is given by some ray $\mathbf{x}+\mathbb{R}\mathbf{1}$, we note that there is a unique representative of this equivalence class whose coordinates sum to zero. In taking such representatives, we can identify the tropical projective torus $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ and the hyperplane $\mathcal{H}=\{\mathbf{x}\in\mathbb{R}^{N}:\sum x_{i}=0\}$. Throughout this paper, we use this homeomorphism to visualize the tropical projective torus (e.g., Figures 1 and 2).
The tropical projective torus is the ambient space containing the tropical Grassmannian, which is equivalent to the space of phylogenetic trees (speyer2004grassmannian). This has sparked the study of tropical data science, formalizing statistical techniques which respect the tropical geometry of the state space (monod2018StatPerspective; yoshida2020science).
The generalized Hilbert projective metric defined above is widely accepted and preferred for the development of geometric statistical tools on the tropical projective torus (yoshida2020science). However, recent work has shown theoretical advantages to the use of an asymmetric metric (comuaneci2023transportation).
@@
Figure 1: The $d_{\mathrm{tr}}$, $d_{\triangle_{\min}}$ and $d_{\triangle_{\max}}$ tropical unit balls in $\mathbb{R}^{3}/\mathbb{R}\mathbf{1}\cong\mathcal{H}=\{\sum x_{i}=0\}$. The solid lines show coordinate directions.
@@
$\mathcal{H}_{\mathbf{a}}=\{\mathbf{x}{\,\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}}:\exists\,i\neq j\text{ s.t. }a_{i}+x_{i}=a_{j}+x_{j}=\max_{k}{(}a_{k}+x_{k}{)}\}.$
@@
Definition 8 (Tropical Convex Hull (develin2004tropical)).
A set $S\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ is max-tropically convex if it contains $\alpha\odot\mathbf{x}\boxplus\beta\odot\mathbf{y}$ for all $\mathbf{x},\mathbf{y}\in S$ and $\alpha,\beta\in\mathbb{R}$. For a finite subset $X=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{{K}}\}\subset\mathbb{R}^{{N}}/\mathbb{R}\mathbf{1}$, the max-plus tropical convex hull of $X$ is the set of all max-tropical linear combinations of points in $X$,
@@ 2.2 Tropical Location Problems
Tropical location problems are a new avenue of research, motivated by the optimization of statistical loss functions; they involve the minimization over some dataset of some loss function which heuristically increases with distance from some kernel. Here we outline the relevant definitions of tropical location problems and their theoretical behavior as presented by comuaneci2024location. We introduce the problem of tropical linear regression (akian2023tropical) as motivation.
@@ 2.2.1 Motivating Example
The problem of tropical linear regression looks to find the best-fit tropical hyperplane which minimizes the maximal distance to a set of data points. While this is an inherently statistical optimization problem, it has recently been proven to be polynomial-time equivalent to solving mean payoff games (akian2023tropical).
A tropical hyperplane is uniquely defined by a cone point $\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, allowing us to formulate the best-fit tropical hyperplane as an optimization problem over ${\mathbf{t}\in\,}\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$.
Definition 9 (Tropical Linear Regression (akian2023tropical)).
@@
$\min_{\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}}f(\mathbf{t})=\max_{k\leq K}d_{\mathrm{tr}}(\mathbf{x}_{k},\mathcal{H}_{\mathbf{t}})=\max_{k\leq K}{\left(\textstyle\max_{i}(\mathbf{x}_{k}-\mathbf{t})_{i}-\text{2ndmax}_{i}(\mathbf{x}_{k}-\mathbf{t})_{i}\right)}$
where $\text{2ndmax}_{i}y_{i}$ denotes some $y_{i}$ such that $\exists\,j\neq i$ such that $\forall\,k\neq i,j$: $y_{j}\geq y_{i}\geq y_{k}$.
This expression for the tropical distance between $\mathbf{x}$ and $\mathcal{H}_{\mathbf{t}}$ was first stated by gartner2008tropical, and has been frequently used in the ongoing study of tropical support vector machines (yoshida2023tropical; tang2024tropical).
In Figure 3, we see a heatmap of this objective function over the tropical projective plane for a sample of size 3. We see immediately that the objective is not convex. The linear regression objective has sparse gradient almost everywhere, which produces valleys. These properties are considered unfavorable extreme cases for gradient methods, but are particularly common in tropical optimization problems. While the tropical linear regression problem is not convex, it is quasi-convex along tropical line segments; that is, its sub-level sets are tropically convex.
@@ 2.2.2 Quasi-Convexity of Tropical Functions
In formalizing tropical location problems, the work of comuaneci2024location first considers functions which increase along tropical geodesics from a kernel $\mathbf{v}$; that is, for any point $\mathbf{t}$ on a geodesic from $\mathbf{v}$ to $\mathbf{w}$, we have that $f(\mathbf{t})\leq f(\mathbf{w})$. This property is equivalent to star-convexity of sub-level sets. To formally state this, we define (oriented) geodesic segments and $\triangle_{\min}$-star convexity of sets.
@@
The (oriented) geodesic segment between $\mathbf{a},\mathbf{b}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, is given by $\mathbf{a},\mathbf{b}]_{\triangle_{\min}}\coloneqq\{\mathbf{x}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}:d_{\triangle_{\min}}(\mathbf{a},\mathbf{x})+d_{\triangle_{\min}}(\mathbf{x},\mathbf{b})=d_{\triangle_{\min}}(\mathbf{a},\mathbf{b})\}$.
@@
A set $S\subseteq\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ is a $\triangle_{\min}$-star-convex set with kernel $\mathbf{v}$ if, for every $\mathbf{w}\in S$, we have $\mathbf{v},\mathbf{w}]_{\triangle_{\min}}\subseteq S$.
Using the above definitions, we can characterize a loss function which is increasing along tropical geodesics as a function with $\triangle_{\min}$-star-convex sub-level sets. The following definition gives a closed form expression for such functions, while Theorem 13 proves their equivalence.
Definition 12 ($\triangle_{\min}$-Star-Quasi-Convex Functions (comuaneci2024location)).
@@
Theorem 13 (Theorem 17 of comuaneci2024location).
@@
These $\triangle_{\min}$-star-convex functions serve as a loss function with respect to a single data point. For example, the tropical metric $d_{\mathrm{tr}}$ is both $\triangle_{\min}$ and $\triangle_{\max}$-star-quasi-convex. The distance to a hyperplane $d_{\mathrm{tr}}(\mathbf{x},\mathcal{H}_{\mathbf{t}})$ is $\triangle_{\max}$-star-quasi-convex in $\mathbf{x}$, but $\triangle_{\min}$-star-quasi-convex in $\mathbf{t}$.
@@
Proposition 15 (Proposition 8 of comuaneci2024location).
Any $\triangle_{\min}$-star-convex set is $\min$-tropically convex. Hence, any $\triangle_{\min}$-star-quasi-convex function is tropically quasi-convex in that it has tropically convex sub-level sets.
@@ 2.2.3 Tropical Location Problems
The $\triangle_{\min}$-star-convex functions defined above are generally a measure of closeness to the kernel $\mathbf{v}$, but in minimizing a statistical loss function we consider proximity to a sample of points. A location problem (laporte2019introduction) measures the closeness to a sample rather than a single kernel.
Definition 16 (Min(/max)-Tropical Location Problem (comuaneci2024location)).
Consider a dataset $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$. For $k\leq K$, let $h_{k}$ be a $\triangle_{\min}$-star-quasi-convex function with kernel $\mathbf{x}_{k}$, and let $g:\mathbb{R}^{K}\rightarrow\mathbb{R}$ be increasing in every coordinate. A min-tropical location problem is the minimization of an objective function $f:\mathbb{R}^{N}/\mathbb{R}\mathbf{1}\rightarrow\mathbb{R}$ given by $f=g(h_{1},\dots,h_{K})$. A max-tropical location problem is defined similarly, where the $h_{k}$ are $\triangle_{\max}$-star-quasi-convex.
The main result of comuaneci2024location is the following: tropical location problems contain a minimum in the tropical convex hull of its dataset $X$.
Theorem 17 (Theorem 20 of comuaneci2024location).
@@
We can now revisit our motivating example (Definition 9), showing that tropical linear regression is not just a min-tropical location problem, but also demonstrates tropical quasi-convexity in that it has tropically convex sub-level sets.
@@
The tropical linear regression problem is a $\min$-tropical location problem with tropically convex sub-level sets.
@@
$f(\mathbf{t})$ $={\max_{k\leq K}\left(\textstyle\max_{i}(\mathbf{x}_{k}-\mathbf{t})_{i}-\text{2ndmax}_{i}(\mathbf{x}_{k}-\mathbf{t})_{i}\right)}$
$=\max_{k\leq K}\text{2ndmin}_{j}(\max_{i}(x_{ki}-t_{i})-x_{kj}+t_{j})$
$=\max_{k\leq K}\text{2ndmin}_{j}(t_{j}-x_{kj}-\min_{i}(t_{i}-x_{ki}{))}$
@@
$\gamma(\mathbf{t})$ $=\text{2ndmin}_{i}t_{i}.$
@@
As the $h_{k}$ are $\triangle_{\min}$-star-convex, their sub-level sets are min-tropically convex by Proposition 15. This is preserved when taking a maximum over samples, so $f$ also has $\min$-tropically convex sub-level sets. ∎